OpenStudy (anonymous):

What value is a discontinuity of x squared plus 2 x plus 3, all over x squared minus x minus 12? A x = -1 B x = -2 C x = -3 D x = -4

OpenStudy (anonymous):

@joyraheb help? or are you busy?

OpenStudy (anonymous):

The discontinuity will be where the denominator equal to 0, which is not allowed

OpenStudy (anonymous):

So just solve x^2-x-12=0 and find those points

OpenStudy (anonymous):

how do I solve it ? I am so stupid with this stuff.

OpenStudy (anonymous):

This is your function: f(x) = (x^2 + 2x + 3)/(x^2-x-12) right?

OpenStudy (anonymous):

I thought the answer was C

OpenStudy (anonymous):

we are not supposed to have anything divided by 0, so let's find these points which make the denominator equal to 0

OpenStudy (anonymous):

x^2-x-12 = 0, that is a quadratic equation, do you know how to find its roots?

OpenStudy (anonymous):

A little bit

OpenStudy (anonymous):

Was I right about it being C. -3?

OpenStudy (anonymous):

there is two points with discontinuity, since we have a quadratic equation. And yes, C is the correct option

OpenStudy (anonymous):

the other point of discontinuit is x = 4

OpenStudy (anonymous):

Let me tell you how to solve a quadratic equation

OpenStudy (anonymous):

yeah that's what I thought because 4 wasn't a choice so it had to be C // but i'll give you a medal thank you so much!

OpenStudy (anonymous):

Consider a general quadratic equation of form: y = a*x^2 + b*x + c

OpenStudy (anonymous):

the roots will be: \[x = \frac{ -b \pm \sqrt{b^2 -4ac} }{ 2a }\]

OpenStudy (anonymous):

;)

OpenStudy (anonymous):

that's how you solve them?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

for example in your situation x^2-x-12 = 0 Here a = 1, b = -1 and c = -12 Just plug in the formula and find the two roots